A spring is attched to the bottom of an empty swimming pool, with axis ot the spring onented vertically. An 80 kg block of wood with density of kgm is xed to the top of spring and water ompletey the The spring is now observed to be streched twice s much as it had been compressed. hollow. a) Determine the the percentage of the blocks total volume that is Ignor any air hollow space. (density of water d 1000 kgm b) If the spring compressed 0.1 m find the spring constant, k
Answer #1

First, volume displaced:
ρV = m ==> ρ/m = V
Where ρ is density, V is volume, m is mass

The block pushes down with a force equal to its weight, 8kgf. Since force =-kx for a spring, when the pool is filled, it lifts with a force of 16kgf against the spring. So, it must displace 92kg of water, because it must lift 16kgf for the spring force, and 16kgf for its own weight.

ρ∙V = m ==> ρ/m = V
total water V = ρ/m = 1000kg/m³ / 92kg = 10.868m³

wood V = ρ/m = 850kg/m³ / 8kg = 106.25m³

void V = ΔV = 106.25m³ - 10.868m³ = 85.38m³

% hollow = Vvoid/ Vtotal = 0.0452m³ / 0.075m³ = 60.3%

I use gravimetric units [kgf], which is not strictly allowed for SI; use mg = W instead of kgf.